# Which Optimizer Works Best for Physics-Informed Neural Networks and Kolmogorov-Arnold Networks?

### v2

@@ Abstract
Physics-Informed Neural Networks (PINNs) have revolutionized the computation of PDE solutions by integrating partial differential equations (PDEs) into the neural network’s training process as soft constraints, becoming an important component of the scientific machine learning (SciML) ecosystem. More recently, physics-informed Kolmogorv-Arnold networks (PIKANs) have also shown to be effective and comparable in accuracy with PINNs. In their current implementation, both PINNs and PIKANs are mainly optimized using first-order methods like Adam, as well as quasi-Newton methods such as BFGS and its low-memory variant, L-BFGS. However, these optimizers often struggle with highly non-linear and non-convex loss landscapes, leading to challenges such as slow convergence, local minima entrapment, and (non)degenerate saddle points. In this study, we investigate the performance of Self-Scaled BFGS (SSBFGS), Self-Scaled Broyden (SSBroyden) methods and other advanced quasi-Newton schemes, including BFGS and L-BFGS with different line search strategies approaches. These methods dynamically rescale updates based on historical gradient information, thus enhancing training efficiency and accuracy. We systematically compare these optimizers -using both PINNs and PIKANs- on key challenging linear, stiff, multi-scale and non-linear PDEs, including the Burgers, Allen-Cahn, Kuramoto-Sivashinsky, and Ginzburg-Landau equations. Our findings provide state-of-the-art results with orders-of-magnitude accuracy improvements without the use of adaptive weights or any other enhancements typically employed in PINNs. More broadly, our results reveal insights into the effectiveness of second-order optimization strategies in significantly improving the convergence and accurate generalization of PINNs and PIKANs.
@@ 1 Introduction
This section begins by providing the necessary background and motivation for our study, highlighting the key challenges and objectives. Following this, we present an overview of the BFGS (Broyden–Fletcher–Goldfarb–Shanno), SSBFGS (Self-Scaled BFGS), and SSBroyden (Self-Scaled Broyden) optimization methods, discussing their theoretical foundations and main characteristics. Section 2 presents a comprehensive study comparing different optimization methods with various line search and trust-region strategies. Specifically, in Subsections 2.1.1 and 2.1.2, we discuss a wide range of scenarios for the Burgers equation. We extend our study to more complex PDEs, including the Allen-Cahn equation in Section 2.3, the Kuramoto-Sivashinsky equation in Subsection 2.4, and the Ginzburg-Landau equation in Section 2.5. The conclusions of our study are summarized in Section 3. Finally, in Appendix A A, we review the performance of optimizers for the Lorenz system. Additionally, in Appendix B LABEL:appendix:Quantitative_Comparison, we provide a quantitative analysis of BFGS and SSBroyden in minimizing the multi-dimensional Rosenbrock function, serving as a pedagogical example to illustrate the impact of dimensionality on optimization performance.
@@ 1.1 Background and Motivation
Currently, the standard optimization algorithms used in PINNs are Adam and quasi-Newton methods, such as BFGS or its low-memory variant, L-BFGS. BFGS and L-BFGS enhance significantly the performance of PINNs by incorporating second-order information of the trainable variable space to precondition it. Although these two optimizers are very commonly employed not only in PINNs but also in other optimization problems because of their theoretical and numerical properties, recent experience in PINNs has shown that other algorithms could outperform them for a great variety of problems [51, 52, 53]. Among all of them we can highlight the Self-Scaled Broyden algorithms [54], which update the approximation to the inverse Hessian matrix using a self-scaling technique. This advanced optimization technique enhances convergence by adaptively scaling gradient updates based on historical error information. Furthermore, [55] investigates self-scaling within quasi-Newton methods from the Broyden family, introducing innovative scaling schemes and updates. Urban et al. [53] further demonstrated that these advanced optimizers, when combined with rescaled loss functions, significantly enhance the efficiency and accuracy of PINNs, enabling smaller networks to solve problems with significantly greater accuracy. As Physics-Informed Neural Networks (PINNs) often suffer from conflicting gradients between multiple loss components (e.g., data and physics losses), resolving these conflicts is essential for stable and accurate training. Recent work by Wang et al. [56] rigorously proves that second-order optimizers, particularly Newton-type methods like SOAP, exhibit positive gradient alignment and can effectively mitigate such conflicts, leading to improved convergence in multi-task optimization settings like PINNs.
@@
To provide an intuitive explanation, Figure 1 illustrates the loss landscape $\mathcal{L}=\mathcal{L}_{D}+\lambda\mathcal{L}_{PDE}$ for the viscous Burgers Equation (16), using the method proposed by [59]. The left subfigure of Figure 1 corresponds to $\lambda=0$, considering only the initial and boundary conditions. This loss landscape is smooth and exhibits well-defined convexity. However, it fits only the initial and boundary condition data and fails to converge to the solution within the domain. In contrast, the right subfigure depicts the loss landscape for $\lambda=1.0$. Unlike the smooth landscape on the left, this landscape is crinkled with numerous local minima and saddle points, including potentially degenerate and non-degenerate saddle points [60]. Consequently, the presence of multiple local minima can trap the optimizer, making it extremely challenging to minimize the loss function effectively. Motivated by this, in this work, we propose and analyze a suite of second-order optimizers and their applicability to linear, nonlinear, stiff, and chaotic PDEs whose solutions are computed by PINN and PIKAN types of approximation. Furthermore, we introduce a set of optimizers tailored to various classes of PDEs, achieving unprecedented convergence to machine precision with both single and double precision arithmetic. To the best of our knowledge, the current study is the most systematic and comprehensive investigation of how optimzers affect the peformance of both PINNs and PIKAns.
@@ 2.5 Ginzburg-Landau equation
The solution domain is defined as $(t,x)\in[0,1]\times[-1,1]^{2}$. Periodic boundary conditions are enforced using a hard-enforcement method with polynomials defined in (18) for $n=4$, applied to each spatial variable $x$ and $y$. Initial conditions, on the other hand, are applied using a soft-enforcement method. Consequently, the loss function integrates contributions from the PDE residuals and the initial conditions. The temporal domain is divided into 5 time windows, each of length $\Delta t=0.2$, with a separate PINN trained for each window. A fully connected neural network is designed to predict the solution fields $u$ and $v$. The network architecture consists of five dense layers, each with 30 neurons, where all hidden layers utilize the hyperbolic tangent ($\tanh$) activation function. The output layer contains two neurons, providing simultaneous predictions for $u$ and $v$. To enhance training efficiency, the model incorporates the adaptive sampling strategy (RAD) algorithm.
@@ 3 Summary
In this study, we conducted a comprehensive comparison of multiple second-order optimizers for PINNs and PIKANs, focusing on BFGS, SSBFGS, SSBroyden, and L-BFGS with Wolfe line-search conditions, as well as BFGS with Backtracking line search and trust-region methods. Utilizing the optax and optimistix library in JAX, we systematically evaluated these optimization techniques across a range of PDEs. We note that we obtained state-of-the-art results by only employing the best optimizers without fine-tuning our loss functions with self-adaptive or attention-based weights or any other enhancemts. We simply pursued a straightforward application of good optimization solvers in order to demonstrate that the big bottleneck for PINNs or PIKANs is the optimization error.
@@
Our initial investigation focused on the Burgers equation, where we evaluated the impact of combining first-order optimizers such as Adam with second-order optimizers such as BFGS and SSBroyden. We then extended our analysis to more challenging PDEs, including the Allen-Cahn equation, the Kuramoto-Sivashinsky equation, and the Ginzburg-Landau equation, providing a detailed evaluation of efficiency and accuracy. The results demonstrate that advanced quasi-Newton methods, particularly SSBroyden, SSBFGS and BFGS with Wolfe line-search, significantly improve the convergence rate and accuracy of PINNs, especially for complex and stiff PDEs. Across all cases, SSBroyden consistently outperformed SSBFGS and BFGS, achieving faster convergence and exhibiting greater robustness in handling complex optimization landscapes. These findings highlight the effectiveness of quasi-Newton methods in accelerating the training of PINNs and enhancing numerical stability. Furthermore, we extended our analysis to PIKANs, replacing traditional multilayer perceptron architectures with KANs utilizing Chebyshev polynomials. In this setting, SSBroyden continued to demonstrate superior optimization performance, reinforcing its robustness across diverse network architectures. Additionally, our study emphasized the importance of implementing PINNs and PIKANs with double-precision arithmetic, which improves numerical stability and enhances optimization efficiency across all tested scenarios.
@@
In Appendix B, we provide a detailed comparison of the number of iterations required to minimize the Rosenbrock function using various optimizers. This analysis highlights how increasing the dimensionality affects the convergence behavior of second-order methods such as BFGS and SSBroyden. We observed that these methods remain efficient across dimensions, while first-order optimizers like ADAM require significantly more iterations and fail to achieve comparable accuracy. The errors reported in this study show that using SSBroyden with the Wolfe line search leads to significantly lower relative $\mathcal{L}_{2}$ errors in all benchmark problems tested compared to the state-of-the-art SOAP method [56]. For instance, in the Burgers equation, our approach reduces the error from $8.06\times 10^{-6}$ to $4.19\times 10^{-8}$. In the Kuramoto–Sivashinsky equation, we achieve an error of $2.65\times 10^{-5}$ over the full time interval $t\in[0,1]$, whereas SOAP reports $3.86\times 10^{-2}$ over a shorter interval $t\in[0,0.8]$. Similar improvements are observed for the Allen–Cahn and Ginzburg–Landau equations (see Table 16).
@@
Our study demonstrates that SSBFGS and SSBroyden with Wolfe line-search are effective and reliable optimizers for training PINNs and PIKANs. Their capacity to handle complex PDEs and their strong convergence properties make them well-suited for advanced applications in scientific machine learning applications. Future research should focus on incorporating domain-specific adaptations to further improve the accuracy, efficiency, and generalization of PINNs and PIKANs, particularly in high-dimensional and multi-scale problem settings.

### v4

@@ Abstract
Physics-Informed Neural Networks (PINNs) have revolutionized the computation of PDE solutions by integrating partial differential equations (PDEs) into the neural network’s training process as soft constraints, becoming an important component of the scientific machine learning (SciML) ecosystem. More recently, physics-informed Kolmogorv-Arnold networks (PIKANs) have also shown to be effective and comparable in accuracy with PINNs. In their current implementation, both PINNs and PIKANs are mainly optimized using first-order methods like Adam, as well as quasi-Newton methods such as BFGS and its low-memory variant, L-BFGS. However, these optimizers often struggle with highly non-linear and non-convex loss landscapes, leading to challenges such as slow convergence, local minima entrapment, and (non)degenerate saddle points. In this study, we investigate the performance of Self-Scaled BFGS (SSBFGS), Self-Scaled Broyden (SSBroyden) methods and other advanced quasi-Newton schemes, including BFGS and L-BFGS with different line search strategies. These methods dynamically rescale updates based on historical gradient information, thus enhancing training efficiency and accuracy. We systematically compare these optimizers – using both PINNs and PIKANs – on key challenging PDEs, including the Burgers, Allen-Cahn, Kuramoto-Sivashinsky, Ginzburg-Landau, and Stokes equations. Additionally, we evaluate the performance of SSBFGS and SSBroyden for Deep Operator Network (DeepONet) architectures, demonstrating their effectiveness for data-driven operator learning. Our findings provide state-of-the-art results with orders-of-magnitude accuracy improvements without the use of adaptive weights or any other enhancements typically employed in PINNs. More broadly, our results reveal insights into the effectiveness of second-order optimization strategies in significantly improving the convergence and accurate generalization of PINNs and PIKANs.
@@ 1 Introduction
This section begins by providing the necessary background and motivation for our study, highlighting the key challenges and objectives. Following this, we present an overview of the BFGS (Broyden–Fletcher–Goldfarb–Shanno), SSBFGS (Self-Scaled BFGS), and SSBroyden (Self-Scaled Broyden) optimization methods, discussing their theoretical foundations and main characteristics. Section 2 presents a comprehensive study comparing different optimization methods with various line search and trust-region strategies. Specifically, in Subsections 2.1.1 and 2.1.2, we discuss a wide range of scenarios for the Burgers equation. We extend our study to more complex PDEs, including the Allen–Cahn equation in Section 2.3, the Kuramoto–Sivashinsky equation in Subsection 2.4, and the Ginzburg–Landau equation in Subsection 2.5. To demonstrate the efficacy of the proposed optimizers, we also solve the Stokes equation for viscous flow in a lid-driven wedge in Subsection 2.6. Furthermore, the performance of SSBFGS and SSBroyden is investigated for data-driven operator learning using Deep Operator Network (DeepONet) in Section 3. The conclusions of our study are summarized in Section 4. Finally, in Appendix A, we review the performance of the optimizers on the Lorenz system. In Appendix B, we provide a quantitative analysis of BFGS and SSBroyden on the multi-dimensional Rosenbrock function, serving as a pedagogical example to illustrate the effect of dimensionality on optimization performance.
@@ 1.1 Background and Motivation
Currently, the standard optimization algorithms used in PINNs are Adam and quasi-Newton methods, such as BFGS or its low-memory variant, L-BFGS. BFGS and L-BFGS enhance significantly the performance of PINNs by incorporating second-order information of the trainable variable space to precondition it. Although these two optimizers are very commonly employed not only in PINNs but also in other optimization problems because of their theoretical and numerical properties, recent experience in PINNs has shown that other algorithms could outperform them for a great variety of problems [51, 52, 53]. Among all of them we can highlight the Self-Scaled Broyden algorithms [54], which update the approximation to the inverse Hessian matrix using a self-scaling technique. This advanced optimization technique enhances convergence by adaptively scaling gradient updates based on historical error information. Furthermore, [55] investigates self-scaling within quasi-Newton methods from the Broyden family, introducing innovative scaling schemes and updates. Urban et al. [53] further demonstrated that these advanced optimizers, when combined with rescaled loss functions, significantly enhance the efficiency and accuracy of PINNs, enabling smaller networks to solve problems with significantly greater accuracy. As PINNs often suffer from conflicting gradients between multiple loss components (e.g., data and physics losses), resolving these conflicts is essential for stable and accurate training. Recent work by Wang et al. [56] demonstrates that second-order optimizers, particularly Newton-type methods like SOAP, exhibit positive gradient alignment and can effectively mitigate such conflicts, leading to improved convergence in multi-task optimization settings, like PINNs. Second-order optimization is also discussed in [57], where the authors present an implementation of a second-order optimizer and demonstrate improvements in both step time and wall-clock time across multiple large-scale tasks spanning diverse domains.
@@
To provide an intuitive explanation, Figure 1 illustrates the loss landscape $\mathcal{L}=\mathcal{L}_{D}+\lambda\mathcal{L}_{PDE}$ for the viscous Burgers Equation (16), using the method proposed by [60]. The left subfigure of Figure 1 corresponds to $\lambda=0$, considering only the initial and boundary conditions. This loss landscape is smooth and exhibits well-defined convexity. However, it fits only the initial and boundary condition data and fails to converge to the solution within the domain. In contrast, the right subfigure depicts the loss landscape for $\lambda=1.0$. Unlike the smooth landscape on the left, this landscape is crinkled with numerous local minima and saddle points, including potentially degenerate and non-degenerate saddle points [61]. Consequently, the presence of multiple local minima can trap the optimizer, making it extremely challenging to minimize the loss function effectively. Motivated by this, in this work, we propose and analyze a suite of second-order optimizers and their applicability to linear, nonlinear, stiff, and chaotic PDEs whose solutions are computed by PINN and PIKAN types of approximation. Furthermore, we introduce a set of optimizers tailored to various classes of PDEs, achieving unprecedented convergence to machine precision with both single and double precision arithmetic. To the best of our knowledge, the current study is the most systematic and comprehensive investigation of how optimizers affect the performance of both PINNs and PIKANs.
@@ 2.5 Ginzburg-Landau equation
The solution domain is defined as $(t,x)\in[0,1]\times[-1,1]^{2}$. Periodic boundary conditions are enforced using a hard-enforcement method with polynomials defined in Equation (18) for $n=4$, applied to each spatial variable $x$ and $y$. Initial conditions, on the other hand, are applied using a soft-enforcement method. Consequently, the loss function integrates contributions from the PDE residuals and the initial conditions. The temporal domain is divided into 5 time windows, each of length $\Delta t=0.2$, with a separate PINN trained for each window. A fully connected neural network is designed to predict the solution fields $u$ and $v$. The network architecture consists of five dense layers, each with 30 neurons, where all hidden layers utilize the hyperbolic tangent ($\tanh$) activation function. The output layer contains two neurons, providing simultaneous predictions for $u$ and $v$. To enhance training efficiency, the model incorporates the adaptive sampling strategy (RAD) algorithm.
@@
Figure 26: Double-precision PINNs for the Stokes equation: The convergence of Physics-Informed Neural Networks (PINNs) for the lid-driven wedge flow governed by the Stokes Equation 23 is studied using the Adam optimizer and SSBFGS, combined with trust-region and backtracking line search methods. All computations are performed using 64-bit floating-point arithmetic. The convergence results indicate that combining SSBFGS with the trust-region line search method significantly outperforms the other two approaches.
#Item Optimizer[# Iters.], Line-search [#Iters.] Relative $l_{2}$ error (u, v, p) Total parameters
1 Adam (18400) + SS-BFGS with Trustregion (419965) ($3.25\times 10^{-4}$, $1.57\times 10^{-4}$, $1.1\times 10^{-1}$) 33,667
2 Adam (18400) + SS-BFGS with backtracking (178741) ($3.26\times 10^{-4}$, $1.60\times 10^{-4}$, $1.1\times 10^{-1}$) 33,667
3 Adam (500000) ($2.7\times 10^{-2}$, $2.58\times 10^{-2}$, $1.1\times 10^{-1}$) 33,667
Table 16: Double-precision PINNs for the Stokes equation: Comparison of relative $l_{2}$ errors, and total number of parameters for PINNs using SSBFGS (with trustregion and backtracking line search) and Adam optimizes. we train the model for all the three cases using same number of model parameters. The absolute and relative tolerance for convergence of SSBFGS optimizers is set $10^{-14}$ and $10^{-14}$.The computation is performed on Nvidia-GPU Card H100 with persistence memory usage of 75% (Total 80 GB) and compute usage of 99%, i.e. 25.35 TFLOPs of total theoretical peak of 25.61 TFLOPs.
2.6 Lid-driven wedge flow
To show the efficacy of proposed optimizer, we finally solve the Stokes equation for a viscous flow in a lid-driven wedge. The Stokes equation is given by
[…]
The boundary conditions and domain configuration are illustrated in Figure 27. This problem is primarily considered to assess the effectiveness of the proposed optimizer for solving PDEs in scenarios involving extremely small values. Moffatt [90] derived an asymptotic solution to Equation 23 using a similarity approach, demonstrating that the strength of eddies depends on the wedge angle. For the specific case depicted in Figure 27, which features a wedge angle of 25.53, the strength of each successive eddy should asymptotically be about 400 times weaker than the preceding one. Accurately resolving such weak eddies requires a highly precise and robust optimizer. This problem has not yet been addressed using the PINNs presented in previous studies, primarily due to two key challenges: the limitations of first-order optimizers and the difficulty of efficiently performing double-precision computations. Even in conventional numerical method these weaker eddies are captured with high-order numerical methods [91]. In this work, we conduct computational experiments to overcome these challenges by comparing our implementation of SSBFGS employing both backtracking and trustregion line search methods—with the commonly used Adam optimizer. Therefore, this example also answers the long standing question of PINN recovering the solution to machine precision. The results of this comparison are discussed in the following paragraphs.
In Figure 27, we show the fluid domain and boundary condition on each edge of the domain with non zero transverse velocity $(u)$ at the top edge and rigid wall (no-slip) condition on the other two edges. To validate the solution obtained by PINNs we compute the reference solution by solving Equation 23 using spectral element method as proposed in [91] and use the Nektar++ codebase [92] to perform the simulation. We show the True streamlines in Figure 27b obtained from the Nektar++. To configure the PINN across all cases, we employ a neural network architecture consisting of 8 fully connected layers, each with 64 neurons, and use the $\tanh$ activation function. A total of 120,000 residual points are sampled uniformly from the domain using $\mathcal{U}[0,1)$. To enhance convergence speed and accuracy, we adopt residual-based attention as proposed in [32], along with the hard imposition of boundary conditions following [93]. Training is conducted for three distinct cases. In the first case, the Adam optimizer is used. In the second and third cases, we employ the SSBFGS optimizer in combination with trust-region and backtracking line search strategies, respectively. For the second-order optimization in these cases, the training is initially warmed up with Adam for 18,401 iterations before switching to SSBFGS. The convergence history for all three cases is presented in Figure 26. As shown in Figure 26, the combination of SSBFGS with the trust-region line search demonstrates superior performance compared to both SSBFGS with backtracking and the Adam optimizer. Specifically, SSBFGS with trust-region converges to a loss value of $10^{-12}$, while SSBFGS with backtracking and the Adam optimizer converge to $10^{-10}$ and $10^{-5}$, respectively.
Training details for all three cases are summarized in Table 16. To evaluate the accuracy of each optimizer, we visualize the resulting streamlines in Figure 28. As illustrated (Figure 28a), the combination of SSBFGS with the trust-region line search successfully recovers all four eddies, including the weakest one with a velocity magnitude of $|u|=10^{-8}$. In contrast, SSBFGS with backtracking (Figure 28b) captures only the dominant eddy accurately, failing to resolve the weaker structures. The streamlines produced using the Adam optimizer (Figure 28c) do not accurately capture any eddies. It is worth noting that although both SSBFGS variants yield similar relative $L_{2}$ errors, this metric reflects the mean error and does not necessarily indicate uniform convergence. Consequently, the streamlines generated by SSBFGS with backtracking significantly deviate, as streamlines are highly sensitive to the precision of the velocity components $(u,v)$. Finally, in Figure 28d, we present the Moffatt eddy strength [90], which compares the centerline transverse velocity magnitude $(|u|)$—indicated by the dashed line in Figure 27(b)—as extracted from Figure 28a, Figure 28b, and Figure 28c. It is evident that the SSBFGS optimizer with trust-region line search most accurately captures the weakest eddies, outperforming both the backtracking variant and the Adam optimizer.
(a) Problem setup for Equation 23
(b) Reference solution
Figure 27: Problem setup and flow condition used for solving Equation 23. Subfigure (a) shows a fluid domain of wedge shape with an aspect ratio of 2:1 and an angle of 25.53. The transeverse velocity $u$ at top lid shows reqgualrized lid and other two boundaries are rigid wall with $u=v=0$. To validate the solution obtained from PINNs, we solve the Equation 23 using the spectral element method [91] and present the streamlines in (b).
(a) SS-BFGS Trustregion
(b) SS-BFGS backtracking
(c) Adam optimizer
(d) Adam optimizer
Figure 28: Solutions of Equation 23 obtained using PINN optimized with (a) SSBFGS optimizer with trustregian based linesearch and subfigure (b) represents the streamlines obtained by pairing the SSBFGS with backtracking linesearch. (c) represents the streamlines obtained while using Adm optimizer. Subfigure (d) shows comaprison of centre-line (shown with dashed line in Figure 27(b) transverse velocity $(|u|)$ in (a), (b) and (c) also known as Moffatt eddy strength.
3 Optimizing Data-Driven DeepONet with Self-Scaled Optimizers
(a) Comparison between predicted and true displacement $u$ using SSBroyden.
(b) Comparison between predicted and true damage $\alpha$ using SSBroyden.
Figure 29: Double-Precision DeepONet Comparison of predicted and true physical fields using the SSBroyden optimizer: (a) displacement $u$, and (b) damage $\alpha$.
In this section, we extend our study to evaluate the performance of second-order optimizers BFGS, SSBFGS, and SSBroyden for DeepONet in a purely data-driven setting. Deep Operator Networks (DeepONets) are neural architectures designed to learn operators—mappings between function spaces—and are particularly effective for solving parametric PDEs [94]. The model consists of two subnetworks, a branch network that encodes input functions sampled at discrete sensor points, and a trunk network that encodes the spatial coordinates at which the output is evaluated. The final output is obtained via an inner product between the branch and trunk outputs, approximating the operator mapping $\mathcal{G}:F\mapsto G$. For full architectural and theoretical details, see [94].
We consider a one-dimensional homogeneous bar under increasing applied displacement to study crack nucleation in a simplified setting. Full details of the dataset are available in our previous work [95]. The dataset consists of phase field $\alpha$ and displacement field $u$, sampled at 101 equispaced spatial nodes over 50 applied displacements $U_{t}\in[0.01,0.5]$. The first 45 displacement steps are used for training, and the remaining 5 for testing. A DeepONet is employed, where the branch network encodes the applied displacements while the trunk network encodes the spatial coordinates. The branch network has 2 hidden layers and the trunk network has 3 hidden layers, each with 20 neurons.
(a) Training loss and error over iterations.
(b) Relative error in damage $\alpha$.
(c) Relative error in displacement $u$.
Figure 30: Double-Precision DeepONet on the 1D dataset. (a) Loss evolution as well as relative $l_{2}$ error over optimizer iterations. (b) Relative $l_{2}$ error in damage variable $\alpha$. (c) Relative $l_{2}$ error in displacement field $u$. The table summarizes training performance for different optimization strategies.
ID Optimizer [Adam, BFGS-Type] Relative $l_{2}$ error (u, $\alpha$) Training Time (s)
1 Adam [15000] $2.1\times 10^{-2}$, $6.2\times 10^{-2}$ 89
2 Adam [10] + BFGS with Wolfe [6000] $5.2\times 10^{-2}$, $8.6\times 10^{-2}$ 399
2 Adam [10] + SS-BFGS with Wolfe [6000] $1.9\times 10^{-2}$, $1.7\times 10^{-2}$ 456
2 Adam [10] + SSBroyden with Wolfe [6000] $5.7\times 10^{-3}$, $6.9\times 10^{-3}$ 444
3 Adam [100] + BFGS with Wolfe [6000] $3.3\times 10^{-2}$, $4.8\times 10^{-2}$ 256
3 Adam [100] + SS-BFGS with Wolfe [6000] $2.9\times 10^{-2}$, $4.8\times 10^{-2}$ 255
3 Adam [100] + SSBroyden with Wolfe [6000] $5.0\times 10^{-3}$, $8.2\times 10^{-3}$ 565
Table 17: Double-Precision DeepONet on the 1D dataset: Comparison of optimizer configurations for the 1D fracture problem. Relative $l_{2}$ errors are reported for displacement $u$ and damage field $\alpha$, and total training time is shown in seconds.
Figure 29 compares (a) the predicted damage field $\alpha$ with the true $\alpha$, and (b) the predicted displacement field $u$ with the true $u$, for 5 testing displacement steps. Figure 30 presents (a) the training loss for four optimizers—Adam, BFGS, SSBFGS, and SSBroyden—where the loss function includes contributions from both the damage field $\alpha$ and the displacement field $u$. It also shows the relative error during training over iterations, demonstrating that SSBroyden and SSBFGS achieve a relative error of $10^{-3}$ with fewer iterations compared to Adam and BFGS. Subfigures (b) and (c) show the absolute error in predicting the damage field $\alpha$ and displacement field $u$, respectively, over the spatial domain. Notably, the error from SSBroyden and SSBFGS remains below 0.0025 for $\alpha$ and around 0.005 for $u$, especially near $x=1$, where the phase field transitions sharply from 0 to 1. In contrast, other optimizers exhibit increased errors at this critical region, highlighting the superior stability and accuracy of SSBroyden. Table 8 summarizes the relative errors and training times for several optimization strategies applied to the same network architecture. In all cases, the network size and data remain fixed. Case 1 uses Adam for 15000 iterations. Cases 2 and 3 combine a warm-up phase using Adam with 6,000 iterations of second-order optimizers. Specifically, Case 2 uses only 10 iterations of Adam before switching to BFGS, SS-BFGS, or SSBroyden, while Case 3 uses a longer warm-up of 100 iterations. The results show that second-order methods, particularly SSBroyden, consistently yield lower errors. While SSBroyden requires slightly longer training times, it achieves the best accuracy across both Case 2 and Case 3 configurations.
@@ 4 Summary
In this study, we conducted a comprehensive comparison of multiple second-order optimizers for PINNs and PIKANs, focusing on BFGS, SSBFGS, SSBroyden, and L-BFGS with Wolfe line-search conditions, as well as BFGS with Backtracking line search and trust-region methods. Utilizing the optax and optimistix library in JAX, we systematically evaluated these optimization techniques across a range of PDEs. We note that we obtained state-of-the-art results by only employing the best optimizers without fine-tuning our loss functions with self-adaptive or attention-based weights or any other enhancements. We simply pursued a straightforward application of good optimization solvers in order to demonstrate that the big bottleneck for PINNs or PIKANs is the optimization error.
@@
Our initial investigation focused on the Burgers equation, where we evaluated the impact of combining first-order optimizers such as Adam with second-order optimizers such as BFGS and SSBroyden. We then extended our analysis to more challenging PDEs, including the Allen-Cahn equation, the Kuramoto-Sivashinsky equation, and the Ginzburg-Landau equation, providing a detailed evaluation of efficiency and accuracy. The results demonstrate that advanced quasi-Newton methods, particularly SSBroyden, SSBFGS and BFGS with Wolfe line-search, significantly improve the convergence rate and accuracy of PINNs, especially for complex and stiff PDEs. Across all cases, SSBroyden consistently outperformed SSBFGS and BFGS, achieving faster convergence and exhibiting greater robustness in handling complex optimization landscapes. These findings highlight the effectiveness of quasi-Newton methods in accelerating the training of PINNs and enhancing numerical stability. Furthermore, we extended our analysis to PIKANs, replacing traditional multilayer perceptron architectures with KANs utilizing Chebyshev polynomials. In this setting, SSBroyden continued to demonstrate superior optimization performance, reinforcing its robustness across diverse network architectures. Additionally, our study emphasized the importance of implementing PINNs and PIKANs with double-precision arithmetic, which improves numerical stability and enhances optimization efficiency across all tested scenarios. Moreover, we evaluated the performance of second-order optimizers —BFGS, SSBFGS, and SSBroyden— for DeepONet in a purely data-driven setting. Specifically, we studied crack nucleation in a one-dimensional homogeneous bar subjected to increasing prescribed displacements.
@@
In Appendix B, we provide a detailed comparison of the number of iterations required to minimize the Rosenbrock function using various optimizers. This analysis highlights how increasing the dimensionality affects the convergence behavior of second-order methods such as BFGS and SSBroyden. We observed that these methods remain efficient across dimensions, while first-order optimizers like ADAM require significantly more iterations and fail to achieve comparable accuracy. The errors reported in this study show that using SSBroyden with the Wolfe line search leads to significantly lower relative $l_{2}$ errors in all benchmark problems tested compared to the state-of-the-art SOAP method [56]. For instance, in the Burgers equation, our approach reduces the error from $8.06\times 10^{-6}$ to $4.19\times 10^{-8}$. In the Kuramoto–Sivashinsky equation, we achieve an error of $2.65\times 10^{-5}$ over the full time interval $t\in[0,1]$, whereas SOAP reports $3.86\times 10^{-2}$ over a shorter interval $t\in[0,0.8]$. Similar improvements are observed for the Allen–Cahn and Ginzburg–Landau equations (see Table 18). A head-to-head comparison of SSBroyden shows that, unlike other second-order optimizers, SSBroyden achieves comparable accuracy with dramatically smaller networks—using just five hidden layers of 30 neurons for the Kuramoto-Sivashinsky equation and three hidden layers of 30 neurons for the Allen–Cahn equation—resulting in substantially reduced training time.
A key question that remains open is the applicability of the proposed family of BFGS-based optimizers for batch training in very large-scale systems. In subsection 2.6, we demonstrated the use of the SSBFGS optimizer to train PINNs with approximately 36,000 parameters, employing 120,000 residual points. However, to fairly compare with first-order optimizers, which efficiently support batch training and thus enable the use of larger neural network architectures, it is important to consider the limitations of BFGS-based methods in such settings. In batch training, the mini-batch changes at each iteration, potentially causing instability in BFGS methods. This is because these methods depend on gradient differences to update Hessian approximations, and the variability in data across batches can introduce inconsistency in those gradients. To address this, [96] proposed the use of overlapping mini-batches for the L-BFGS optimizer to ensure some continuity between iterations and improve the stability of gradient difference estimates. We implemented this multi-batch strategy for the viscous Burgers equation and achieved comparable accuracy to that obtained using full-batch training. Nevertheless, a comprehensive evaluation of multi-batch training for all variants of BFGS optimizers lies beyond the scope of the present study. Such an investigation would require a detailed analysis of how noisy gradient estimates, arising from batch sampling, affect the Hessian approximation, and how the condition number of the Hessian varies with the type of PDE being solved. This is left for future work.
@@
In summary, our study demonstrates that SSBFGS and SSBroyden with Wolfe line-search are effective and reliable optimizers for training PINNs and PIKANs. Their capacity to handle complex PDEs and their strong convergence properties make them well-suited for advanced applications in scientific machine learning applications. Future research should focus on incorporating domain-specific adaptations to further improve the accuracy, efficiency, and generalization of PINNs and PIKANs, particularly in high-dimensional and multi-scale problem settings.